# The value of hedging against energy storage uncertainties when designing energy parks

### v1

@@ 3.4 Stochastic Programming for system design
Stochastic Programming is used to optimize the capacities of wind and solar generation and energy storage in the energy park system. Specifically, linear scenario programming is used due to its computational efficiency, enabling the use of an hourly resolved model with multiple scenarios, and its resulting prevalent use in the literature for designing district- and national-scale energy systems [42, 43, 44, 45].
@@
The design optimization aims to minimize the expected annualized operating cost of the energy park, the average total cost of running the energy park to provide the industrial load over all scenarios considered. The total cost is made up of the capital costs of the wind, solar, and energy storage, the cost (or profit) of energy bought from/sold to the grid, and the cost of carbon emissions associated with purchased grid electricity. The optimization is subject to energy $[\,\cdot\,]^{+}$ and $[\,\cdot\,]^{-}$ represent the positive and negative parts of the argument respectively. (Eq. 7a), storage capacity (Eq. 7b & Eq. 7c), and grid capacity (Eq. 7h) constraints for operation within each scenario $m$. As well as a budget constraint on total capital expenditure (Eq. 7j). During operation, electricity can be bought from and sold to the electricity grid (Eq. 7f), and the renewable generation can be dynamically curtailed (Eq. 7g).
Scenarios are generated by sampling from the distributions of renewable generation, industrial load, and storage technology performance described in Sec. 3.3. The sampled storage capacity cost and lifetime values are combined to produce annualized capacity cost samples for each technology, $p^{s}_{i,m}$. The number of scenarios that can be considered during optimization is limited by high computational cost. Therefore, scenario reduction [30, 46, 47] is used to improve the statistical representation of possible scenarios in the optimization. A large initial sample of scenarios is drawn, the optimization is performed considering each scenario separately, and a subset of scenarios which best represent the distribution of individually optimized costs is selected for use in the Stochastic Program.
@@
${\min}$ $\qquad p^{\textrm{w}}C^{\textrm{w}}+p^{\textrm{pv}}C^{\textrm{pv}}+{{\sum}}_{m}\>p_{m}\left({\sum}_{i}\>p^{s}_{i,m}C^{s}_{i}+{\sum}_{t}\left(p^{e}[t]E^{\text{grid}}_{m}[t]+\,p^{c}\,c[t]\left[E^{\text{grid}}_{m}[t]\right]^{+}\right)\right)$ (7)
@@
subject to $\qquad\textrm{SoC}_{i,m}[t{+}1]=\textrm{SoC}_{i,m}[t]+\sqrt{\eta_{i}}\left[E_{i,m}[t]\right]^{+}-1/\sqrt{\eta_{i}}\left[\,E_{i,m}[t]\right]^{-}$ (7a)
@@
$\qquad\textrm{SoC}_{i,m}[0]=\textrm{SoC}_{i}^{0}$ (7e)
@@
$\eta_{i}$ Round-trip efficiency of storage technology $i$
$p^{s}_{i,m}$/kWh/yr Annualized capacity cost of storage technology $i$ in scenario $m$
@@
$\textrm{SoC}_{i}^{0}$ kWh Initial state-of-charge of storage technology $i$
$p^{\text{w}}$/kWp/yr Annualized capacity cost of offshore wind
$p^{\text{pv}}$/kWp/yr Annualized capacity cost of solar PV
$p^{e}[t]$/kWh Price of grid electricity at time $t$
$p^{c}$/kgCO2 Nominal carbon price
@@
$B^{\text{cap}}$ Capital budget constraint
@@ 3.4.1 Accounting for risk aversion
The optimization objective in Eq. 7 assumes that the energy park designer is risk-neutral and is only interested in minimizing the average operating cost of the system, as is standard in the energy system design literature. However, when investing billions of Euros in large energy infrastructure projects, energy firms are typically risk-averse in their decision making, and are concerned about the risk of high operating costs. Previous studies have investigated the impact of risk aversion on energy system design [30, 31, 32], but none have considered how risk aversion affects the value of uncertainty reduction or optionality during decision making.
To account for the impact of risk aversion during energy park design, a risk measure is included in the objective function. The Conditional Value-at-Risk (CVaR) is used as it is a coherent risk measure compatible with Linear Programming [48]. For a linear scenario program where the cost and probability of each scenario are $c_{m}$ and $p_{m}$ respectively, an optimization minimizing the CVaR is given by,
$\min$ $\quad\xi+\frac{1}{\alpha}\sum\limits_{m}p_{m}\eta_{m}$ (8)
@@
When optimized, this objective is equal to the expected value of the costs in the $\alpha$ right-tail of scenarios, $\mathbb{E}_{m:\,c_{m}>\xi_{\alpha}}\{c_{m}\}$, i.e. the average of the highest $\alpha\%$ of costs across the scenarios.
As VoI and VoO are differences in average costs, to allow valid comparison between the risk-neutral and risk-averse cases, we develop a risk-averse objective which maintains the scale of the cost. This objective is equivalent to a weighted average of the scenario costs, where the costs in the $\alpha\%$ right-tail are weighted by a factor $n{+}1$ relative to others. So, risk aversion is represented by the designer ‘caring’ about high operating costs more than others, meaning the objective is still a physical cost, and so can be validly compared across cases. This risk-averse objective is,
@@
where $c_{m}$ is the objective value in each scenario as specified in Eq. 7, and both the original constraints for Eq. 7 and the constraints from Eq. 8 are imposed. When optimized this objective becomes $\left(\mathbb{E}\{c_{m}\}+n\alpha\,\mathbb{E}_{m:\,c_{m}>\xi_{\alpha}}\{c_{m}\}\right)/(1+n\alpha)$.
@@ 4 Results & Discussion
4.1 Design without optionality
[…]
Storage technology Total operating cost (€m/yr) Carbon emissions (ktCO2/yr) Storage capacity (GWh) Cost of storage (€m/yr)
@@ 4.1 Design without optionality
[…]
The average total cost of the final energy park design over the samples is €102.0m/yr, whereas the cost of the initial design was €124.1m/yr. Therefore, the Value of Information (VoI) associated with designing the energy park after R&D has been performed to reduce uncertainty in the performance of NaS storage is €22.1m/yr (18% of initial design cost). This cost reduction comes from the system designer’s improved ability to trade-off wind generation and storage capacity depending on the relative cost of energy arbitrage, and better make use of the capital budget. In some cases where the cost of NaS storage is low and its efficiency is high, the storage capacity in the final design is more than double that initial chosen. So, assuming a 20 year project lifetime, having the option to adjust the capacities of wind and storage procured from their suppliers after R&D is worth €442m to the energy park developer. As a result, they should be willing to pay more (per unit capacity) to a supplier offering a flexible contract.
@@
4.2 Design with optionality
@@
The average cost of the energy park designed with storage technology optionality is €88.7m/yr, compared to €102.0m/yr if only NaS is available for final design. Therefore, the value of retaining storage optionality for the final design of the energy park (VoO) is €13.3m/yr (13% of cost without optionality). With a 20yr system lifetime, this optionality in storage technology selection is worth €266m to the energy park developer. Long-duration energy storage demonstrator projects funded by the UK government have received around €1m for proof-of-concept systems [49, 50] and €10m for mid-scale demonstrators [51, 52]. The cost of developing and contracting Li-ion, VRFB, and CAES storage is conservatively estimated at €20m each. Therefore, hedging bets on storage technology performance and maintaining optionality in technology choice through the energy park design process provides a net reduction in average cost of €206m (10%) to the energy park developer.
@@
4.3 Sensitivity to measurement model
A key assumption made in the probabilistic model of uncertain storage performance in Sec. 3.3 was the level of uncertainty reduction provided by performing R&D and measuring the demonstrator system. This is encapsulated by the uncertainty reduction factor parameter $r$ in Eq. 6, which quantifies how closely the performance of the demonstrator matches that of a large-scale system, and so how much uncertainty remains after measurement. Finding data from real-world storage systems to estimate this parameter is extremely challenging, as actual performance and cost data are highly commercially sensitive, and few large-scale systems currently exist.
Sensitivity analysis is performed to determine whether the conclusions provided by the Value of Optionality analysis remain valid. The process from Sections 4.1 & 4.2 is repeated for values of $r$ over the feasible range. Table 5 shows VoI and VoO values for each case. For both VoI and VoO, the greater the uncertainty reduction provided by measurement (lower $r$), the greater the benefit to design. This is because lower uncertainty allows final designs with less hedging to be used, which achieve lower cost. While there will be some threshold of $r$ above which the VoO is less than the cost of R&D, even with the lowest level of uncertainty reduction tested, $r=0.5$, retaining storage technology selection optionality is still economically worthwhile. So the conclusions of the base case analysis are robust to the level of uncertainty reduction. The variation of VoI with $r$ is larger than that of VoO, as VoO compares differences in designs both made with reduced uncertainty, so both costs in the difference benefit (reduce) from lower uncertainty.
@@ 4.3 Sensitivity to measurement model
Table 5: Sensitivity of VoI and Value of Optionality to level of uncertainty reduction.
@@ 4.4 Use of multiple storage technologies
So far it has been assumed that only a single energy storage technology is installed in the energy park to perform energy arbitrage. The different characteristics of the energy storage technologies (such as efficiency, cost per energy capacity, and discharge ratio) make them suited for arbitrage over different time scales [7]. As trends in renewable generation and grid electricity price also occur over a range of time scales, the profit from energy arbitrage could be increased by installing multiple energy storage technologies, and sizing them for the volume of the arbitrage they are best suited to.
The investigation is repeated for the case where two storage technologies are used in the energy park. This means two technologies are selected and developed following the initial design. Fig. 10 plots an example of the operation of the energy park with two storage technologies, for the initial system design with Li-ion and CAES. The state-of-charge traces for Li-ion (orange) and CAES (dark blue) demonstrate the different arbitrage behaviours of the two technologies. The Li-ion operates at short time scales with small traded energy volumes, while the CAES operates over longer time scales with greater energy volumes.
@@
For the energy park with two technologies, the average cost of the final designs is €87.5m/yr for the case without optionality, and €85.7m/yr for the case with optionality. Table 6 compares these costs to the results for the system with only one storage technology from Sections 4.1 & 4.2. While using two storage technologies provides a substantial cost reduction for the case without optionality in the final design, €14.5m/yr (14%), with optionality this cost reduction falls to €3.0m/yr (3%). However, this cost saving is still significant (€60m over a 20yr system lifetime), and shows that multiple energy storage technologies should be used for arbitrage in energy parks.
Design case Total operating cost (€m/yr) VoI/VoO (€m/yr) Carbon emissions (ktCO2/yr) Cost of storage (€m/yr)
@@
The VoO is now €1.8m/yr (2.1%), or €36m over the 20yr lifetime. This is less than the estimated €40m cost of developing the remaining two storage technologies not selected in the initial design. So, when two technologies are used, it is not worthwhile hedging bets against uncertain storage performance by retaining optionality. This means selecting the best two storage technologies is robust to the performance uncertainty. The VoI however remains high, €26.4m/yr, meaning that contract flexibility and adapting the final design are still very valuable.

### v4

@@ 3.4 Stochastic Programming for system design
Stochastic Programming is used to optimize the capacities of wind and solar generation and energy storage in the energy park system. Specifically, linear scenario programming is used due to its computational efficiency, enabling the use of an hourly resolved model with multiple scenarios, and its resulting prevalent use in the literature for designing district- and national-scale energy systems, see [37] & [48].
@@
The design optimization aims to minimize the expected annualized cost of the energy park, the average total cost of running the energy park to provide the industrial load over all scenarios considered. The total cost is made up of the capital costs of the wind, solar, and energy storage, the cost (or profit) of energy bought from/sold to the grid, and the cost of carbon emissions associated with purchased grid electricity. The optimization is subject to energy conservation444[]+[$\cdot\,$$^{+}$ and $\cdot\,$$^{-}$ represent the positive and negative parts of the argument respectively. (Eq. 7a), storage capacity (Eq. 7b & Eq. 7c), and grid capacity (Eq. 7h) constraints for operation within each scenario $m$. As well as a budget constraint on total capital expenditure (Eq. 7j). During operation, electricity can be bought from and sold to the electricity grid (Eq. 7f), and the renewable generation can be dynamically curtailed (Eq. 7g).
Scenarios are generated by sampling from the distributions of renewable generation, industrial load, and storage technology performance described in Sec. 3.2 & 3.3. The sampled storage capacity cost and lifetime values are combined to produce annualized capacity cost samples for each technology, $p^{s}_{i,m}$. The number of scenarios that can be considered during optimization is limited by high computational cost. Therefore, scenario reduction [38] is used [^5] Specifically the fast-forward scenario reduction algorithm from [18], implemented by [17]. to improve the statistical representation of possible scenarios in the optimization. A large initial sample of scenarios is drawn, the optimization is performed considering each scenario separately, and a subset of scenarios which best represent the distribution of individually optimized costs is selected for use in the Stochastic Program.
@@
${\min}$ $\qquad p^{\textrm{w}}C^{\textrm{w}}+p^{\textrm{pv}}C^{\textrm{pv}}+{{\sum}}_{m}\>\rho_{m}\left({\sum}_{i}\>p^{s}_{i,m}C^{s}_{i}+{\sum}_{t}\left(p^{e}[t]E^{\text{grid}}_{m}[t]+\,p^{c}\,c[t]\left[E^{\text{grid}}_{m}[t]\right]^{+}\right)\right)$ (7)
@@
subject to $\qquad\textrm{SoC}_{i,m}[t{+}1]=\textrm{SoC}_{i,m}[t]+\sqrt{\eta_{i,m}}\left[E_{i,m}[t]\right]^{+}-1/\sqrt{\eta_{i,m}}\left[\,E_{i,m}[t]\right]^{-}$ (7a)
@@
$\qquad\textrm{SoC}_{i,m}[0]=\textrm{SoC}^{0}C^{s}_{i}$ (7e)
@@
$\eta_{i,m}$ Round-trip efficiency of storage technology $i$ in scenario $m$
$p^{s}_{i,m}$ €/kWh/yr Annualized capacity cost of storage technology $i$ in scenario $m$
@@
$\textrm{SoC}^{0}$ Initial state-of-charge of storage (fraction of capacity)
$p^{\text{w}}$ €/kWp/yr Annualized capacity cost of offshore wind
$p^{\text{pv}}$ €/kWp/yr Annualized capacity cost of solar PV
$p^{e}[t]$ €/kWh Price of grid electricity at time $t$
$p^{c}$ €/kgCO2 Nominal carbon price
@@
$B^{\text{cap}}$ €/yr Annualized capital budget constraint
@@ 3.4.1 Accounting for risk aversion
The optimization objective in Eq. 7 assumes that the energy park designer is risk-neutral and is only interested in minimizing the average operating cost of the system, as is standard in the energy system design literature. However, when investing billions of Euros in large energy infrastructure projects, energy firms are typically risk-averse in their decision making, and are concerned about the risk of high operating costs. Previous studies have investigated the impact of risk aversion on energy system design, such as [38] & [45], but none have considered how risk aversion affects the value of uncertainty reduction or optionality during decision making.
To account for the impact of risk aversion during energy park design, a risk measure is included in the objective function. The Conditional Value-at-Risk (CVaR) is used as it is a coherent risk measure compatible with Linear Programming [41]. For a linear scenario program where the cost and probability of each scenario are $c_{m}$ and $\rho_{m}$ respectively, an optimization minimizing the CVaR is given by,
$\min$ $\quad\xi+\frac{1}{\alpha}\sum\limits_{m}\rho_{m}\eta_{m}$ (8)
@@
When optimized, this objective is equal to the expected value of the costs in the $\alpha$ right-tail of scenarios, $\mathbb{E}_{m\mathrel{\mathop{\ordinarycolon}}\,c_{m}>\xi_{\alpha}}\{c_{m}\}$, i.e. the average of the highest $\alpha\%$ of costs across the scenarios.
As VoI and VoO are differences in average costs, to allow valid comparison between the risk-neutral and risk-averse cases, we develop a risk-averse objective which maintains the scale of the cost. This objective is equivalent to a weighted average of the scenario costs, where the costs in the $\alpha\%$ right-tail are weighted by a factor n+1n{+}1 relative to others. So, risk aversion is represented by the designer ‘caring’ about high operating costs more than others, meaning the objective is still a physical cost, and so can be validly compared across cases. This risk-averse objective is,
@@
where $c_{m}$ is the objective value in each scenario as specified in Eq. 7, and both the original constraints for Eq. 7 and the constraints from Eq. 8 are imposed. When optimized this objective becomes $\left(\mathbb{E}\{c_{m}\}+n\alpha\,\mathbb{E}_{m\mathrel{\mathop{\ordinarycolon}}\,c_{m}>\xi_{\alpha}}\{c_{m}\}\right)/(1+n\alpha)$.
@@ 4 Results & Discussion
Initial designs for the energy park using each storage technology are produced by applying the Stochastic Program (SP) model (described in Sec. 3.4), using the prior distributions of renewable generation, industrial load, and storage technology performance (defined in Sec. 3.2 & 3.3). Table 4 compares the performance of these designs. Sodium-Sulphur high-temperature batteries (NaS) are found to provide the lowest total system cost, so this design is selected, and NaS is chosen for procurement and R&D. The design sizes $441\text{\,}\mathrm{M}\mathrm{W}$ wind and $500\text{\,}\mathrm{M}\mathrm{W}$ solar generation capacity, with $1.51\text{\,}\mathrm{G}\mathrm{W}\mathrm{h}$ of NaS storage to support the system. The average total cost of this system is €124.1m/yr, with €10.7m/yr expected to be spent on batteries. If no energy storage were used in the energy park, $486\text{\,}\mathrm{M}\mathrm{W}$ of wind and $500\text{\,}\mathrm{M}\mathrm{W}$ of solar generation would be installed, leading to an average cost of €183.7m/yr. Therefore, installing bulk energy storage for arbitrage reduces the average cost of the energy park system by €59.6m/yr (32%), and reduces the average operational carbon emissions by 20.9 ktCO2/yr (20%).
Storage technology Total cost (€m/yr) Carbon emissions (ktCO2/yr) Storage capacity (GWh) Cost of storage (€m/yr)
@@
4.1 Design updating without optionality & the Value of Information
To investigate the final design of the energy park given the uncertainty in the performance of NaS storage, 250 possible measurement values obtained from the demonstrator system built during R&D were sampled for each performance parameter using the probabilistic model of storage performance (see Table 2). For each sampled measurement, the SP model is applied using the corresponding posterior distribution of large-scale system performance to produce a final design. Fig. 6 plots the wind generation and NaS storage capacities of the final designs for each sampled measurement, and compares them to the initial design. All designs install the maximum $500\text{\,}\mathrm{M}\mathrm{W}$ of solar.
The average total cost of the final energy park design over the samples is €102.0m/yr, whereas the cost of the initial design was €124.1m/yr. Therefore, the Value of Information (VoI) associated with designing the energy park after R&D has been performed to reduce uncertainty in the performance of NaS storage is €22.1m/yr (18% of initial design cost). This cost reduction comes from the system designer’s improved ability to trade off wind generation and storage capacity depending on the relative cost of energy arbitrage, and better make use of the capital budget. In some cases where the cost of NaS storage is low and its efficiency is high, the storage capacity in the final design is more than double that initially chosen. Assuming a 20 year project lifetime, having the option to update the capacities of wind and storage procured from their suppliers after R&D is worth €442m to the energy park developer. As a result, they should be willing to pay more (per unit capacity) to a supplier offering a flexible contract.
@@ 4.1 Design updating without optionality & the Value of Information
4.2 Designing with optionality & the Value of Optionality
@@
The average cost of the energy park designed with storage technology optionality is €88.7m/yr, compared to €102.0m/yr if only NaS is available for final design. Therefore, the value of retaining storage optionality for the final design of the energy park (VoO) is €13.3m/yr (13% of cost without optionality). With a 20 year system lifetime, this optionality in storage technology selection is worth €266m to the energy park developer. Long-duration energy storage demonstrator projects funded by the UK government have received around €1m for proof-of-concept systems [13] and €10m for mid-scale demonstrators [12]. So, the cost of developing and contracting Li-ion, VRFB, and CAES storage is conservatively estimated at €20m each. Therefore, hedging bets on storage technology performance and maintaining optionality in technology choice through the energy park design process provides a net reduction in average cost of €206m (10%) to the energy park developer.
@@
4.3 Sensitivity analysis
A key assumption made in the probabilistic model of uncertain storage performance in Sec. 3.3 was the level of uncertainty reduction provided by measuring the demonstrator system built during R&D. This is encapsulated by the uncertainty reduction factor parameter $r$ in Eq. 6, which quantifies how closely the designer believes the performance of the demonstrator matches that of a large-scale system, and so how much uncertainty remains after measurement. Finding data from real-world storage systems to estimate this parameter is extremely challenging, as actual performance and cost data are highly commercially sensitive, and few large-scale systems currently exist.
Sensitivity analysis is performed to determine whether the conclusions provided by the Value of Optionality analysis remain valid. The process from Sec. 4.1 & 4.2 is repeated for values of $r$ over the feasible range. Table 5 shows VoI and VoO values for each case. For both VoI and VoO, the greater the uncertainty reduction provided by measurement (lower $r$), the greater the benefit to design. This is because lower uncertainty allows final designs with less hedging to be used, which achieve lower cost. While there will be some threshold of $r$ above which the VoO is less than the cost of R&D, even with the lowest level of uncertainty reduction tested, $r=0.5$, retaining storage technology selection optionality is still economically worthwhile. So the conclusions of the base case analysis are robust to the level of uncertainty reduction. The variation of VoI with $r$ is larger than that of VoO, as VoO compares differences in designs both made with reduced uncertainty, so both costs in the difference benefit (reduce) from lower uncertainty.
@@ 4.3 Sensitivity analysis
Table 5: Sensitivity of Value of Information and Value of Optionality to level of uncertainty reduction.
During operation the energy park is able to trade with the grid to reduce the cost of the energy it supplies to the hydrogen plant, but must pay a cost for any associated carbon emissions. However, due to lack of data availability, uncertainty in grid electricity price and carbon emissions were not included in the uncertainty model. So, a similar sensitivity analysis is performed to check whether the pricing and carbon data assumed in the base case are representative, or whether changes in this data have a significant effect on the resulting VoI and VoO values.
Three alternative cases are tested: the 2023 pricing data is increased and decreased by 10% respectively, and pricing and carbon data from 2024 is used. Table 6 presents the VoI and VoO values for each of these cases. While the VoI is found to vary significantly, by up to 75%, in all of the alternative cases it is greater than the base case. This suggests that design flexibility could be even more valuable than indicated by the initial results. As before, the changes in VoO values are smaller than for the VoI. And again, in all cases the VoO is greater than the cost of developing the additional storage technologies, so retaining optionality is still found to be economically worthwhile. Therefore, the conclusions from Sec. 4.1 & 4.2 are not affected by the electricity price and carbon emissions data used.
Case Prior operating cost (€m/yr) Value of Information (€m/yr) Value of Optionality (€m/yr)
2023 prices & carbon (base case) 124.1 22.1 13.3
2023 prices $10\%$ 124.9 39.2 16.1
2023 prices $10\%$ 151.5 32.1 12.1
2024 prices & carbon 80.9 38.2 17.1
Table 6: Sensitivity of Value of Information and Value of Optionality to electricity prices and carbon emissions
Whether or not retaining optionality in storage technology selection is worthwhile is determined by the comparison between the VoO and the cost of retaining optionality. The cost of optionality is therefore a critical driver of this conclusion. Were the €20m cost of building a demonstrator system used in the initial analysis (see Sec. 4.2) to double, this would correspond to an annualized cost of €6m/yr for developing all three additional storage technologies. However, even this doubling of a pessimistic cost estimate is still significantly smaller than any of the VoO values found across the sensitivity analyses, the smallest of which is €9.8m/yr from the extreme case of low uncertainty reduction, $r=0.5$. So, the assumptions made about the cost of optionality do not affect the conclusions drawn from the analysis, as the VoO is sufficiently large to provide strong confidence that storage optionality is worthwhile for the design of the studied energy park.
@@ 4.4 Use of multiple storage technologies
So far it has been assumed that only a single energy storage technology is installed in the energy park to perform arbitrage. The different characteristics of the energy storage technologies (such as efficiency, cost per energy capacity, and discharge ratio) make them suited for arbitrage over different time scales [21]. As trends in renewable generation and grid electricity price also occur over a range of time scales, the profit from energy arbitrage could be increased by installing multiple energy storage technologies, and sizing them for the volume of the arbitrage they are best suited to.
The investigation is repeated for the case where two storage technologies are used in the energy park. This means two technologies are selected and developed following the initial design. Fig. 10 plots an example of the operation of the energy park with two storage technologies, for the initial system design with Li-ion and CAES. The state-of-charge traces for Li-ion (orange) and CAES (dark blue) demonstrate the different arbitrage behaviours of the two technologies. As the Linear Program simulating the energy park (Eq. 7) simultaneously optimizes over all operational variables, it is able to dynamically dispatch the two storage technologies, scheduling them to best suit their characteristics and maximize the combined performance. In this case, best performance is achieved by operating the Li-ion at short time scales with small traded energy volumes, while the CAES is operated over longer time scales with greater energy volumes.
@@
Design case Total cost (€m/yr) VoI/VoO (€m/yr) Carbon emissions (ktCO2/yr) Cost of storage (€m/yr)
@@
For the energy park with two technologies, the average cost of the final designs is €87.5m/yr for the case without optionality, and €85.7m/yr for the case with optionality. Table 7 compares these costs to the results for the system with only one storage technology from Sec. 4.1 & 4.2. While using two storage technologies provides a substantial cost reduction for the case without optionality in the final design, €14.5m/yr (14%), with optionality this cost reduction falls to €3.0m/yr (3%). However, this cost saving is still significant (€60m over a 20yr system lifetime), and shows that multiple energy storage technologies should be used for arbitrage in energy parks.
The VoO is now €1.8m/yr (2.1%), or €36m over the 20 year lifetime. This is less than the estimated €40m cost of developing the remaining two storage technologies not selected in the initial design. So, when two technologies are used, it is not worthwhile hedging bets against uncertain storage performance by retaining optionality. This means selecting the best two storage technologies is robust to the performance uncertainty. The VoI however remains high, €26.4m/yr, meaning that contract flexibility and updating the final system design are still very valuable.